OpenStudy (anonymous):

simplify each expression sqrt45n^5

OpenStudy (saifoo.khan):

\[\sqrt{45} = \sqrt{9 * 5} = 3\sqrt5\]

OpenStudy (anonymous):

Thanks man ur amazing

OpenStudy (saifoo.khan):

\[\sqrt{n^5} = n^2\sqrt n\]Now combine them. \[3n^2 \sqrt{5n}\]

OpenStudy (saifoo.khan):

Welcome. :)

OpenStudy (saifoo.khan):

Mind if i ask you which grade math is that?

OpenStudy (anonymous):

8th

OpenStudy (saifoo.khan):

Cool. ^_^